### v1

@@ A.2 Fair Cuts
Let $f$ denote the flow constructed in the proof of Theorem 5.1. We decompose $f$ into flow paths from $s$ to $t$ and scale down each flow path so that the flow crossing every edge in the fair cut $E_{H}(S,T)$ equals $\frac{2}{\phi}$. Let $f^{\prime}$ denote the resulting flow.
For the first item, we delete from $f^{\prime}$ all flow paths that contain an edge $(v,t)$, where $t$ is the sink and $v\in S=\bar{A^{\prime}}\setminus\bar{A}$. Denote the resulting flow $f^{\prime\prime}$. The requested flow is achieved by restricting the flow paths in $f^{\prime\prime}$ to their subpath that intersects $T\setminus\{t\}$ (that is, deleting the prefix in $S$ and the last vertex $t$).
We proceed to the second item, and prove it in two steps. We first route flow from the new cut edges $E_{G}(A^{\prime},\bar{A^{\prime}})\setminus E_{G}(A,\bar{A})$ to $E_{G}(A,\bar{A})$, and then we route the flow from $E_{G}(A,\bar{A})$ to $\bar{A}$ so that every $v\in\bar{A}$ receives at most $O(\phi\log n\mu(v))$ flow.
For the first step, observe that there is a flow $f_{1}$ of congestion $1.5=O(1)$ in $G[\bar{A^{\prime}}]$ such that every $v\in\bar{A^{\prime}}\setminus\bar{A}=S\setminus\{s\}$ sends $(deg_{G}-deg_{G[\bar{A^{\prime}}]})(v)$ flow and receives at most $1.5(deg_{G}-deg_{G[\bar{A}]})(v)$ flow. This flow is obtained by scaling the flow paths of $f^{\prime}$ by $\phi/2$ and then restricting them to their prefix from $s$ until the first edge in $T$.
In the second step, note that the cut $(A,\bar{A})$ was produced in the cut-matching step. Therefore, by Corollary A.3, there exists a flow $f_{2}$ of congestion $O(1)$ in $G[\bar{A}]$ such that every $v\in\bar{A}$ sends $(deg_{G}-deg_{G[\bar{A}]})(v)$ flow and receives at most $O(\phi\log n\cdot\mu(v))$ flow.
The final flow is as follows: We first route according to $f_{1}$. Each $v\in\bar{A}$ must then send a total of $2.5(deg_{G}-deg_{G[\bar{A^{\prime}}]})(v)$ flow: this consists of $(deg_{G}-deg_{G[\bar{A^{\prime}}]})(v)$ flow required by the lemma and an additional $1.5(deg_{G}-deg_{G[\bar{A^{\prime}}]})(v)$ flow from $f_{1}$. We route this flow using a scaled version of $f_{2}$, which incurs at most $2.5\cdot O(1)$ congestion. ∎
@@ Appendix B Algebraic Tools
Fact B.1 (Exercise IX.3.3 [4]).
@@
Lemma B.2 (In the proof of Theorem IX.3.5 [4]).
@@
Theorem B.3 (Symmetric Rearrangement; Theorem A.2 [9]).
@@ Appendix C Projection Lemmas
Lemma C.1 (Gaussian Behavior of Projections; Variation of Lemma 3.5 [5]).
@@
Lemma C.1 follows from Lemma 3.5 of [5] as follows: For a single projection, condition (2) fails to hold with probability at most $e^{-\alpha\log n/4}=n^{-\alpha/4}$. By the union bound it fails for some vector with probability $n^{1-\alpha/4}$. For this failure probability to be small we need $\alpha$ to be sufficiently large; If $\alpha\geq 8$ then the failure probability is at most $n^{-\alpha/8}$.

### v2

@@ A.2 Fair Cuts
Let $H$ be the auxiliary graph in the proof of Theorem 5.1, and let $f$ denote the flow constructed therein. We decompose $f$ into flow paths from $s$ to $t$ and scale down each flow path so that the flow crossing every edge of $E_{H}(S,T\setminus\{t\})$ equals $\frac{2}{\phi}$ (because $(S,T)$ is a fair cut). Let $f_{1}$ denote the resulting flow. Let $f_{2}\mathrel{\mathop{:}}=\frac{\phi}{2}f_{1}$. Note that $f_{2}$ routes exactly one unit of flow across every edge of $E_{H}(S,T\setminus\{t\})$, at most $\frac{3}{2}$ units of flow across the remaining edges of $H\setminus\{t\}$, and at most $\frac{\phi}{2}\mu(v)$ units across each edge $(v,t)$.
For the first item, we delete from $f_{2}$ all flow paths that contain an edge $(v,t)$, where $t$ is the sink and $v\in S=\{s\}\cup\overline{A^{\prime}}\setminus\overline{A}$. Denote the resulting flow by $f_{3}$. The desired flow is then obtained by restricting the flow paths in $f_{3}$ to their suffix, starting from the edge crossing the fair-cut $(S,T)$.
We proceed to the second item, and prove it in two steps. We first route flow from the new cut edges $E_{G}(A^{\prime},\overline{A^{\prime}})\setminus E_{G}(A,\overline{A})$ to $E_{G}(A,\overline{A})$, and then we route the flow from $E_{G}(A,\overline{A})$ to $\overline{A}$ so that every $v\in\overline{A}$ receives at most $O(\phi\log n\cdot\mu(v))$ units of flow.
For the first step, observe that there is a flow $f^{(1)}$ of congestion $\frac{3}{2}=O(1)$ in $G[\overline{A^{\prime}}]$ such that every $v\in\overline{A^{\prime}}$ sends $\left(\deg_{G}-\deg_{G[\overline{A^{\prime}}]}\right)\!(v)$ units of flow and each $u\in\overline{A}$ receives at most $\frac{3}{2}\left(\deg_{G}-\deg_{G[\overline{A}]}\right)\!(u)$ units of flow. This flow is obtained by restricting the flow paths of $f_{3}$ to their prefix from $s$ until the first edge in the fair-cut $(S,T)$, and then reversing them, so that they route flow from the cut $(S,T)$ towards $s$. In the last edge $(v_{1},s)$ of each flow path, instead of sending the flow to $s$, we will send it to the vertex $v_{0}$ of $G[\overline{A}]$ corresponding to the edge $(v_{0},v_{1})\in E(A,\overline{A})$.
In the second step, note that the cut $(A,\overline{A})$ was produced in the cut-matching step. Therefore, by Corollary A.4, there exists a flow $f^{(2)}$ of congestion $O(1)$ in $G[\overline{A}]$ such that every $v\in\overline{A}$ sends $\left(\deg_{G}-\deg_{G[\overline{A}]}\right)\!(v)$ units of flow and receives at most $O(\phi\log n\cdot\mu(v))$ units of flow.
The final flow is as follows. We first route according to $f^{(1)}$. Each $v\in\overline{A}$ must then send at most $\frac{3}{2}\!\left(\deg_{G}-\deg_{G[\overline{A^{\prime}}]}\right)\!(v)$ units of flow, which are the units it receives from $f^{(1)}$. We route this flow using a scaled version of $f^{(2)}$, specifically $\frac{3}{2}f^{(2)}$, which incurs a congestion of $O(1)$. ∎
@@ Appendix B Algebraic Tools
Fact B.1 ([bhatia2013matrix, Exercise IX.3.3]).
@@
Lemma B.2 (In the proof of [bhatia2013matrix, Theorem IX.3.5]).
@@
Theorem B.3 (Symmetric Rearrangement; [orecchia2008partitioning, Theorem A.2]).
@@ Appendix C Projection Lemmas
Lemma C.1 (Gaussian Behavior of Projections; Variation of [khandekar2009graph, Lemma 3.5]).
@@
Lemma C.1 follows from [khandekar2009graph, Lemma 3.5] as follows: For a single projection, condition (2) fails to hold with probability at most $e^{-\alpha\log n/4}=n^{-\alpha/4}$. By the union bound it fails for some vector with probability $n^{1-\alpha/4}$. For this failure probability to be small we need $\alpha$ to be sufficiently large; If $\alpha\geq 8$ then the failure probability is at most $n^{-\alpha/8}$.